Required information [The following information applies to the questions displayed below.) You have just been hired as a finaLydex Company Comparative Income Statement and Reconciliation This Year Last Year Sales (all on account) $ 15,940,000 $ 14,382. You decide next to assess the companys stock market performance. Assume that Lydexs stock price at the end of this year
Answer #1

Ans:

This year

Last Year

Earning Per Share

= Net Income/No of Shares

= Net Income/No of Shares

=1,120,000/(7500,000/75)

=1,152,900/(7500,000/75)

11.20

11.53

Dividend Payout Ratio

= Divined Per share/ Earning Per Share

= Divined Per share/ Earning Per Share

=3.2/11.20

=5.76/11.53

28.57%

49.96%

Dividend Yield Ratio

= Dividend Per Share/ Market Value Per Share

= Dividend Per Share/ Market Value Per Share

=3.2/110

=5.76/78

=2.9%

=7.38%

Price Earning Ratio

= Market Price Per Share/ Earning Per share

= Market Price Per Share/ Earning Per share

=110/11.20

=78/11.53

9.82

6.76

Book Value Per Share

= Total Shareholder's Equity/ No of Shares

= Total Shareholder's Equity/ No of Shares

=9,880,000/(100000)

=9080,000/(100000)

98.8

=90.80